if a,b,c,d are positive real numbers such that a/3=a+b/4=a+b+c/5=a+b+c+d/6 then
a/b+2c+3d

Grade:

7 Answers

kasilaxmi
9 years ago
h
kasilaxmi
9 years ago
h
kasilaxmi
9 years ago
hi farzana
farzana
9 years ago
hello chandu.....
kasilaxmi
9 years ago
hey post raledu kada
Y RAJYALAKSHMI
9 years ago
Let a/3 = a + b / 4 = a + b + c / 5 = a + b + c + d/ 6 = k
then, a = 3k, b = k, c = k, d = k
Β 
a/b + 2c + 3d = 3k/k + 2k + 3k = 3k/6k = 1/2
Samrat Hadke
2 years ago
Correct option is
A

21​

Β 

Given that,Β 3a​=4a+b​=5a+b+c​=6a+b+c+d​
Β 
So, we haveΒ 3a​=4a+b​
β‡’a=3b
Β 
We have,Β 3a​=5a+b+c​
β‡’a=3c
Β 
We have,Β 3a​=6a+b+c+d​
β‡’a=3d
Β 
So, the value ofΒ b+2c+3da​=3a​+32a​+aa​=2aa​=21​
Β 
Therefore, the correct option isΒ A. Β½
Β